OpenStudy (anonymous):

2. What is a solution of x² + 6x = –5? (1 point) A.x = –6 B.x = –1 C.x = 1 D.x = 6

OpenStudy (anonymous):

B. x = -1

OpenStudy (anonymous):

(-1)^2 + 6(-1) = 1 - 6 = -5

OpenStudy (anonymous):

so is answer B correct?

OpenStudy (anonymous):

actually, there are two solution for x, they're -5 and -1, but since -5 is not found on the options, therefor the answer will be -1

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

(x+1)(x+5) x=-1 x=-5

OpenStudy (anonymous):

and I meant to set that equal to zero. So you get why there are those two solutions for x.